# OpeNLGauge: An Explainable Metric for NLG Evaluation with Open-Weights LLMs

### v1

@@ Appendix H Output Examples
Listing 1: Annotator prompt template for data-to-text task
@@
Listing 2: Supervisor prompt template for merging of annotation
@@
Listing 3: Prompt template for the fine-tuned LLM
@@
G-Eval 0.189 0.351 0.198 0.338 0.269
GPT-3.5 0.304 0.385 0.306 0.242 0.294
GPT-4 0.213 0.405 0.320 0.302 0.359
@@
Prometheus 0.161 0.150 0.173 0.169 0.211
Auto-J 0.179 0.084 0.141 0.176 0.127
TIGERScore 0.160 0.221 0.200 0.215 0.204
InstructScore 0.194 0.300 0.247 0.222 0.273
Themis 0.298 0.395 0.333 0.259 0.380
OpeNLGaugeens 0.234 0.415 0.205 0.341 0.299
- Command R+ 104B 0.236 0.299 0.047 0.210 0.198
- Gemma 2 27B 0.254 0.334 0.275 0.317 0.295
- Llama 3.1 Nemotron 70B 0.178 0.284 0.226 0.281 0.242
- Mistral Large 2 123B 0.129 0.359 0.215 0.311 0.254
- Qwen 2.5 70B 0.226 0.347 0.245 0.315 0.283
@@
Table 14: Segment-level Spearman ($r$) correlations of different metrics on SFRES and SFHOT. Inf. = informativeness, Nat. = naturalness.
@@
Table 15: Segment-level Pearson ($r$) correlations of various metrics on HANNA. Coh. = coherence, Rel. = relevance, Eng. = engagement, Emp. = empathy, Sur. = surprise, Com. = complexity, Avg. = average.
@@
Table 16: Segment-level Spearman correlations ($\rho$) of various metrics on HANNA. Coh. = coherence, Rel. = relevance, Eng. = engagement, Emp. = empathy, Sur. = surprise, Com. = complexity, Avg. = average.
@@
Table 17: Segment-level Kendall ($\tau$) correlations of various metrics on HANNA. Coh. = coherence, Rel. = relevance, Eng. = engagement, Emp. = empathy, Sur. = surprise, Com. = complexity, Avg. = average.
@@
Table 18: Segment-level Pearson ($r$) correlations of different metrics on Wiki-DA.
@@
Data: Adolfo_Suárez_Madrid–Barajas_Airport | operatingOrganisation | ENAIRE
@@
Text: Adolfo Suárez Madrid–Barajas Airport is a major international airport located near the heart of the city, offering a wide range of air services.
@@
Evaluation aspect: Faithfulness – Extent to which the information in the text is supported by thedata Explanation: This phrase is considered an error because the provided data does not mention the airport’s size (major/international), location relative to the city (near the heart), or any information that could imply these details. The text introduces unsupported information, negatively affecting faithfulness. Severity: 4 Explanation: Similar to Error 1, this phrase introduces information not present in the provided data. There is no mention of the types or variety of air services offered by the airport, making this an unsupported claim that affects faithfulness. Severity: 3 Overall Score: Poor (2/5) Explanation of the Score: The generated text contains significant errors related to faithfulness, as it introduces substantial information not supported by the provided data. While the text does not contradict any data points (since it doesn’t directly address them), its lack of adherence to the given facts and introduction of unsubstantiated claims warrant a ‘Poor’ rating. For an ‘Excellent’ score, the text should have strictly adhered to the provided data, which includes specific details about the airport’s operating organisation, runway length, elevation, and runway name, none of which were accurately represented or utilized in the generated text.
@@
Figure 14: Example error span annotation provided by OpeNLGauge for the RDF-to-text task. The text generated from the data is evaluated for faithfulness.
Emotional Context: A is terrified Dialogue History: A: I moved into this new house and was woken up with this horrible screeching noise in the middle of the night. B: Terrifying! What was it? A: It turned out to be the sump pump, I guess it was old. But yes, it scared the daylights out of me! B: Oh good. I bet! I wouldn’t have been able to go back to sleep after that! Do you have to replace the pump? A: I did, yeah. It was actually still in working condition but I couldn’t stand the sound and I’m a light sleeper. No more scary noises at 3am! Response: that ’s great! i’m glad you’re able to deal with it now! Evaluation aspect: Empathy – Extent to which the response shows understanding of the feelings of the person talking about their experience Explanation: The response acknowledges the resolution of the issue but fails to explicitly acknowledge A’s emotional distress (being terrified) caused by the event. Using a casual, positive phrase like "that’s great" might come across as dismissive of the initial fear experienced. Severity: 3 Explanation: While this part of the response shows some form of relief for A’s current situation, it doesn’t directly address or validate A’s past emotional state (terror) related to the incident. It focuses more on the practical outcome rather than offering empathy for the emotional experience. Severity: 2 Overall Score: Fair (3/5) Explanation of the Score: The response demonstrates a basic understanding that the issue is resolved and expresses relief for the person, which shows some level of consideration. However, it lacks direct acknowledgment and validation of A’s terrified emotional state, preventing it from being scored higher in empathy. The errors identified, while present, do not completely negate the empathetic intent, hence a ’Fair’ score rather than ’Poor’ or ’Unacceptable’.
@@
Figure 15: Example error span annotation provided by OpeNLGauge for the dialogue response generation task. The response is evaluated for empathy. In addition to dialogue history, the input also includes the emotional context.
@@
Summary: By A new study that college are more likely to engage in energy - binge drinking when alcohol than non - . Evaluation aspect: Fluency – Formal quality of individual sentences of the summary. A fluent sen-tence should be grammatical, natural and easy to understand. Explanation: The sentence is fragmented and lacks proper structure. It starts with an unnecessary preposition ("By") and contains several grammatical errors, such as missing words and incorrect phrasing. This makes the sentence difficult to understand and unnatural. Severity: 5 Overall Score: Unacceptable (1/5) Explanation of the Score: The summary is very disfluent due to significant grammatical errors and lack of coherence, making it extremely difficult to comprehend.
@@
Figure 16: Example error span annotation provided by OpeNLGauge for the summarization task. The summary is evaluated for fluency. Note that source text is not included as the input for fluency evaluation.

### v3

@@ Appendix I Inter-annotator agreement between ensemble models
Listing 1: Annotator prompt template for the data-to-text task.
@@
Listing 2: Consolidator prompt template for merging of individual annotations.
@@
Listing 3: Prompt template for the fine-tuned OpeNLGaugeft metric.
@@
G-Eval (GPT-4) 0.189 0.351 0.198 0.338 0.269
LLM Evaluation (GPT-3.5) 0.304 0.385 0.242 0.294 0.306
LLM Evaluation (GPT-4) 0.213 0.405 0.302 0.359 0.320
Prometheus 0.161 0.150 0.211 0.169 0.173
Auto-J 0.179 0.084 0.176 0.127 0.141
TIGERScore 0.160 0.221 0.215 0.204 0.200
InstructScore 0.194 0.300 0.222 0.273 0.247
Themis 0.298 0.395 0.259 0.380 0.333
OpeNLGaugeens 0.234 0.415 0.205 0.341 0.299
• Command R+ 104B 0.239 0.298 0.215 0.311 0.266
• Gemma 2 27B 0.254 0.334 0.275 0.317 0.295
• Llama 3.1 Nemotron 70B 0.178 0.284 0.047 0.209 0.179
• Mistral Large 2 123B 0.129 0.359 0.226 0.281 0.249
• Qwen 2.5 72B 0.226 0.347 0.245 0.315 0.283
@@
Table 16: Segment-level Spearman ($\rho$) correlations of different metrics on SFRES and SFHOT. Inf. = informativeness, Nat. = naturalness.
@@
Table 17: Segment-level Pearson ($r$) correlations of different metrics on HANNA. Coh. = coherence, Rel. = relevance, Eng. = engagement, Emp. = empathy, Sur. = surprise, Com. = complexity, Avg. = average.
@@
Table 18: Segment-level Spearman ($\rho$) correlations of different metrics on HANNA. Coh. = coherence, Rel. = relevance, Eng. = engagement, Emp. = empathy, Sur. = surprise, Com. = complexity, Avg. = average.
@@
Table 19: Segment-level Kendall ($\tau$) correlations of different metrics on HANNA. Coh. = coherence, Rel. = relevance, Eng. = engagement, Emp. = empathy, Sur. = surprise, Com. = complexity, Avg. = average.
@@
Table 20: Segment-level Pearson ($r$) correlations of different metrics on Wiki-DA. Meaning = Meaning preservation.
@@
Figure 15: Inter-annotator agreement (Krippendorff’s $\alpha$) between all LLMs in the ensemble for all meta-evaluation datasets. For each dataset, the coefficient is computed over all evaluation aspects.
Figure 16: Pairwise Spearman ($\rho$) correlations of individual LLM scores for all meta-evaluation datasets. For each dataset, the correlations are computed over all evaluation aspects.
Figure 17: Pairwise inter-annotator agreements (Cohen’s $\kappa$) of individual LLM scores for all meta-evaluation datasets. For each dataset, the coefficients are computed over all evaluation aspects.
Figure 18: Distribution of errors detected by the ensemble LLMs in outputs rated with maximum score by human annotators in SummEval. Top: Frequencies of numbers of detected errors per evaluated output. Middle: Frequencies of mean severity levels assigned to detected error per output. Values larger than 0 are binned to ranges $[a,b)$, where $0<a<=5$ and $b=a+1$. Bottom: Frequencies of maximum severity levels assigned to detected errors per output.
Figure 19: Distribution of errors detected by the ensemble LLMs in outputs rated with maximum score by human annotators in TopicalChat. Note that groundedness evaluations are excluded from the analysis, as the dataset contains only binary ratings for this aspect. Top: Frequencies of numbers of detected errors per evaluated output. Middle: Frequencies of mean severity levels assigned to detected error per output. Values larger than 0 are binned to ranges $[a,b)$, where $0<a<=5$ and $b=a+1$. Bottom: Frequencies of maximum severity levels assigned to detected errors per output.
Model/Method QAGS SummEval TopicalChat SFRES/SFHOT HANNA Wiki-DA
Command R+ 104B 0.681 0.394 0.332 0.266 0.323 0.681
Gemma 2 27B 0.643 0.442 0.481 0.295 0.366 0.767
Llama 3.1 Nemotron 70B 0.669 0.451 0.660 0.179 0.393 0.752
Mistral Large 2 123B 0.645 0.502 0.598 0.249 0.376 0.769
Qwen 2.5 72B 0.651 0.478 0.496 0.283 0.301 0.783
Average 0.688 0.533 0.652 0.299 0.409 0.837
Average w/o outliers 0.677 0.538 0.623 0.289 0.411 0.825
Majority vote 0.654 0.504 0.556 0.290 0.381 0.785
Median 0.668 0.509 0.594 0.296 0.401 0.803
Min 0.641 0.467 0.622 0.265 0.389 0.776
Table 21: Segment-level Spearman ($\rho$) correlations of different score aggregation methods. For each dataset, correlations are averaged across aspects. Individual LLMs are included for comparison. Best ensemble results for each dataset are highlighted in bold.
Data:
Adolfo_Suárez_Madrid–Barajas_Airport | operatingOrganisation | ENAIRE
@@
Text: Adolfo Suárez Madrid–Barajas Airport is a major international airport located near the heart of the city[1], offering a wide range of air services[2].
@@
Evaluation aspect:Faithfulness – Extent to which the information in the text is supported by the data [1] Explanation: This phrase is considered an error because the provided data does not mention the airport’s size (major/international), location relative to the city (near the heart), or any information that could imply these details. The text introduces unsupported information, negatively affecting faithfulness. Severity: 4 [2] Explanation: Similar to Error 1, this phrase introduces information not present in the provided data. There is no mention of the types or variety of air services offered by the airport, making this an unsupported claim that affects faithfulness. Severity: 3 Overall score: Poor (2/5) Explanation of the score: The generated text contains significant errors related to faithfulness, as it introduces substantial information not supported by the provided data. While the text does not contradict any data points (since it doesn’t directly address them), its lack of adherence to the given facts and introduction of unsubstantiated claims warrant a ‘Poor’ rating. For an ‘Excellent’ score, the text should have strictly adhered to the provided data, which includes specific details about the airport’s operating organisation, runway length, elevation, and runway name, none of which were accurately represented or utilized in the generated text.
@@
Figure 20: Example error span annotation provided by OpeNLGauge for the RDF-to-text task. The generated text is evaluated for faithfulness with respect to the input data.
Emotional Context: A is guilty Dialogue History:
A: One time I skipped my friend’s birthday party to watch TV
B: I hope something good was on A: There was! But I felt so bad afterwards. I’m a monster Response: It’s okay[1], everyone makes mistakes sometimes. Next time, try to find a balance between social commitments and personal enjoyment[2].!
Evaluation aspect:Empathy – Extent to which the response shows understanding of the feelings of the person talkingabout their experience [1] Explanation: While the intention is to comfort, the phrase "It’s okay" might come across as dismissive of A’s expressed guilt and remorse, potentially undermining their emotional experience. A more empathetic approach could acknowledge the negative feelings before offering reassurance. Severity: 2 [2] Explanation: This part of the response shifts focus from acknowledging and validating A’s current emotional state (guilt) to providing advice for future actions. While helpful, it doesn’t directly address or show understanding of A’s feelings about their past action, slightly lacking in empathy. Severity: 3 Overall score: Fair (3/5) Explanation of the score: The response demonstrates some level of empathy by attempting to reassure and offer a solution, indicating an understanding that mistakes happen. However, it falls short in fully acknowledging and validating the speaker’s emotional experience (guilt), hence not scoring higher than ’Fair’. To achieve a ’Good’ or ’Excellent’ score, the response would need to more directly address and empathize with A’s feelings of guilt and self-perceived monstrosity.
@@
Figure 21: Example error span annotation provided by OpeNLGauge for the dialogue response generation task. The response is evaluated for empathy. In addition to dialogue history, the input also includes the emotional context of the person talking about their experience.
@@
Source: [NOT PROVIDED] Summary: By A new study that college are more likely to engage in energy - binge drinking when alcohol than non - .[1] Evaluation aspect:Fluency – Formal quality of individual sentences of the summary. A fluent sentence should begrammatical, natural and easy to understand. [1] Explanation: The sentence is fragmented and lacks proper structure. It starts with an unnecessary preposition ("By") and contains several grammatical errors, such as missing words and incorrect phrasing. This makes the sentence difficult to understand and unnatural. Severity: 5 Overall score: Unacceptable (1/5) Explanation of the score: The summary is very disfluent due to significant grammatical errors and lack of coherence, making it extremely difficult to comprehend.
@@
Figure 22: Example error span annotation provided by OpeNLGauge for the summarization task. The summary is evaluated for fluency. Note that source text is not included, as it is not relevant for evaluation of fluency.